OpenStudy (anonymous):

What is the value of z in the equation 3z + 5 = 4z − 8? −13 −1 1 13

OpenStudy (midhun.madhu1987):

3z + 5 = 4z − 8 Move all terms containing z to left side and numbers to right side: 3z - 4z = -8 - 5 Simplify to get the answer...

OpenStudy (anonymous):

what

OpenStudy (midhun.madhu1987):

what is 3z - 4z =?

OpenStudy (anonymous):

i dont know what z is

OpenStudy (anonymous):

-13?

OpenStudy (midhun.madhu1987):

3z + 5 = 4z − 8 3z - 4z = -8 - 5 -z = -13 z = 13

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Thank you